OpenStudy (anonymous):

The retina of a human eye can detect light when radiant energy incident on it is at least 4.0 x 10^-17 J. For light of 505 nm wavelength, how many photons does this energy correspond to?

OpenStudy (cuanchi):

you have to calculate the energy of one photon E= h c /wavelength (in meters) convert the 505 nm to meters h= 6.626 x 10^(-34) Js (Plank's constant) c= 3.0 x 10^8 m/s (speed of light) After you calculate the energy of one photon of 505nm you have to divide the total energy 4.0 x 10^-17J by the energy of the one photon to calculate how many photons you have in that amount of energy

OpenStudy (anonymous):

I got 9.75 x 10^-37 but I think I did it wrong.

OpenStudy (aaronq):

you should post your work, but yeah, that number is not right - you should get a number of photons, not energy

OpenStudy (anonymous):

\[E = \frac{ 6.626 x 10^{-34}J/s x 3.0 x 10^{8}m/s }{5.05x10^{-7} m }\]

OpenStudy (anonymous):

\[\frac{ 3.9 x 10^{-19} J }{ 4.0 x 10^{-17}J }\]

OpenStudy (anonymous):

That's how I did it

OpenStudy (aaronq):

good stuff. it's good up until the division. you were supposed to the divide the threshold energy, 4.0*10^-17 J by the energy of the photon you can calculated.

OpenStudy (anonymous):

so the 4.0/3.9?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

1.0256 x 10-36

OpenStudy (aaronq):

you're pluggin it into the calculator incorrectly

OpenStudy (aaronq):

the answer is 102.5641, you should try to get this number

OpenStudy (anonymous):

What order did you put it into your calculator?

OpenStudy (aaronq):

i use this http://web2.0calc.com/

OpenStudy (aaronq):

havent used a normal, handheld calculator in a long time

OpenStudy (anonymous):

Think my calculator is just messing up.

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (aaronq):

yeah, it's tricky pluggin in negative exponents. no problem

OpenStudy (cuanchi):

@staldk3 Try to use parentheses when you enter the denominator with a power of ten, otherwise the calculator will understand that you want to multiply everything by 10^-19 in place to divide by the 10^-19 (4.0 x 10^-17) / (3.9 x 10^-19) = 1.03 x 10^2 Otherwise the calculator understand (4.0 x 10^-17 / 3.9) x 10^-19 = 1.03 x 10^-36